Reggie is 5 years older than Lani. 5 years ago the sum of their ages was 25. How old are they now?
Answer by mananth(16946)

let lani be x years
Reggie = x+5 years
...
5 years ago
lani x-5
Reggie = x+5 -5
..
x-5 +x+5-5 = 25
2x-5 =25
add 5 to both sides
2x-5+5 = 25 +5
2x=30
x=15 lani's age
Reggie's age = x+5 = 15+5=20 years